### v2

# Alpha-Particle Monopole Form Factors with Ab Initio No-Core Shell Model

Abstract
The state-of-the-art ab initio nuclear many-body approaches with modern nuclear forces are challenged by the recent experimental measurement of the monopole form factor of the transition in the particle [Kegel et al., Phys. Rev. Lett. 130, 152502 (2023)]. We investigate the elastic and inelastic transition form factors using the ab initio no-core shell model (NCSM). We observe a good convergence of both form factors with respect to the basis size employing the Daejeon16 nucleon-nucleon () interaction. Our NCSM results are very close to the effective interaction hyperspherical harmonic calculations using plus three-nucleon interactions based on the chiral effective field theory which take into account the continuum effects via the Lorentz integral transform. The significant difference between the ab initio results with various modern nuclear interactions and of some of them with the recent experimental data provides motivations for deeper investigation of this observable.
I Introduction
The particle, i. e., the 4He nucleus, is the lightest doubly magic nucleus and can be addressed as a natural laboratory for inspecting the inter-nucleon interactions and nuclear many-body methods [1]. The 4He ground state is deeply bound with the binding energy per nucleon of about MeV and has been extensively studied by nuclear physicists over the years. Its first excited state has the same spin and parity as the ground state. The state lies about MeV above the H threshold but about MeV below the opening of the He channel [2]. Therefore it is an unstable resonance decaying by proton emission.
Recently, the A1 Collaboration at Mainz Microtron (MAMI) has performed precision measurements of the monopole form factor of the transition via inelastic electron scattering [3]. The new data improves the precision significantly compared to the previous measurements [5, 4, 6].
Several theoretical investigations have been performed for the transition form factor using realistic inter-nucleon interactions. In Ref. [7], a bound state technique, which expands the wave functions in a Gaussian basis, was used to calculate this observable with the simplified Argonne V8’ (AV8’) nucleon-nucleon () interaction [8] plus a phenomenological three-nucleon (3) interaction. The results of Ref. [7] describe relatively well the experimental data. The effective interaction hyperspherical harmonic (EIHH) calculations in Ref. [9] using the more accurate Argonne V18 (AV18) potential [8] plus the Urbana IX (UIX) 3 force [10] and chiral effective field theory (EFT) based interactions ( interaction at N3LO [11] and 3 interaction at N2LO [12]) revealed sizable deviations from the experimental data. Therefore, there is a puzzle: why do modern inter-nucleon interactions fail to explain the MAMI data [3]? As was pointed out in Ref. [13], “the form factor may serve as a tool, or ‘magnifying glass’, for probing small contributions to nuclear forces”.
Following up the MAMI measurement, a calculation within the no-core Gamow shell model in the coupled-channel representation (NCGSM-CC) [14], employing a version of the chiral interaction at N3LO [11], claimed to explain the experimental data. However, the study of Ref. [14] neglected 3 forces and revealed a strong sensitivity to , the position of the resonance with respect to the H threshold. It was found that by fitting the transition form factor to the experimental inelastic form factor through varying instead of calculating within the NCGSM, a good agreement with the Kegel et al. data can be obtained at MeV [14, Erratum] that is much smaller than the experimentally observed value MeV.
In Ref. [15], the ab initio nuclear lattice EFT calculations also reproduced the experimental data, albeit with a simplified nuclear interaction formulated on a periodic cubic box [16] which is not known to provide a high quality description of the phase-shift data. In addition, we note that the monopole transition density presented in [15, Fig. 2] after the Fourier transform results in an about 20 higher form factor with respect to that presented in [15, Fig. 1], which raises a concern over the robustness of the claimed agreement with experiment.
The most recent results with the hyperspherical harmonic (HH) method in Ref. [17] reproduced well the inelastic form factor experimental data by using the chiral interactions ( interaction at N3LO [11] together with 3 interaction at N2LO [18, 19]). However, this calculation results in a form factor that significantly differs from that obtained in Ref. [9] using the same but another parameterization of the chiral 3 force. The differences in the 3 interaction parametrization are not expected to produce an approximately disagreement in the values of the transition form factor around its maximum. This difference may result from the fact that the H channel allowed in Ref. [17] includes both isospins and thus suppressing the role of the component of the resonance, which provides a dominant contribution to the monopole transition, and therefore the transition form factor in the calculations of Ref. [17] is reduced. Interestingly, in Ref. [17] the position of the state is predicted to be about MeV above the H threshold, much below the experimental value.
The issue of describing the inelastic monopole form factor using modern nuclear forces remains unresolved. The differences between various reported theoretical calculations and at least some of them with the MAMI experimental data raise significant questions. Additional studies of this observable could shed further light on the Hamiltonian and method dependences.
In this letter, we present an alternative ab initio calculation for the monopole form factor of the transition in 4He with the no-core shell model (NCSM). We perform the calculation with the Daejeon16 interaction [20]. This interaction is based on the N3LO chiral EFT interaction of Ref. [11], softened via a similarity renormalization group (SRG) transformation so as to improve convergence of ab initio studies, modified off-shell to mimic three-nucleon forces and provide one of the best descriptions of nuclei with [21].
II Theoretical methods
The monopole transition form factor is related to the point-proton transition density by
where is the charge of 4He under consideration [15]. The proton form factor takes care of the proton finite size. As was done in Ref. [15], we use a microscopic from Ref. [22] corresponding to the proton charge radius of fm, which is very close to the most recent experimental data fm [23]. A phenomenological expression ( in fm-1), which corresponds to the proton charge radius of fm, has been frequently used in calculations. We found that the monopole form factors of 4He calculated with the microscopic and phenomenological proton form factors are nearly indistinguishable over the range of the experimental data. The point-proton transition density is the matrix element of the point-proton density operator , where denotes the position vector of th proton, between the ground state and the first excited state [7, 15],
The transition form factor transforms to the elastic form factor if we replace the point-proton transition density in Eq. (1) by the ground state point-proton density. There are various definitions for the monopole transition form factor in available references. In order to compare with the most recent experimental data of Ref. [3] and make comparisons with other theoretical calculations, one should ensure that the corresponding elastic monopole form factors satisfy
We use the ab initio NCSM [24] to calculate the wave functions of the state and the state. The NCSM has been extensively used in studies of light nuclei (see, e. g., Refs. [25, 26, 27, 28, 29, 30, 31, 32, 33]). In the NCSM, the nuclear wave functions are obtained by diagonalizing the chosen nuclear Hamiltonian in a truncated Slater determinant harmonic oscillator (HO) basis characterized by the basis oscillator parameter . The truncation of the model space is determined and labeled by the number of HO excitation quanta, , relative to the minimum number of quanta required by the Pauli principle. We test convergence by showing calculated quantities vs , and we report how these quantities approach their asymptotic values as increases.
In this work, we perform NCSM calculations with the code MFDn using the Lanczos algorithm [34, 35, 36, 37]. We use the Daejeon16 NN interaction [20] and add the Coulomb interaction between the protons. We adopt the Lawson method [38, 39] to ensure that only states which are free from spurious center of mass (CoM) excitation remain in the calculated spectrum. We employ the method in Ref. [40] to obtain the translational invariant density in Eq. (1).
The state of 4He is a resonance sitting just above the 3H threshold. Therefore the continuum state obtained with the NCSM in the above manner is viewed as based on a discretized approximation of the continuum. This approximation becomes more reliable as the basis size increases. For example, in Ref. [30] we calculated the polarizability, an inverse energy weighted sum rule, of 4He by directly solving for thousands of discretized states and reached robust convergence with increasing.
III Results and discussion
We start with calculations of the elastic 4He form factor. In Fig. 1 (a) we present the point-proton density distribution of the 4He ground state for with the same MeV. We observe good exponential tail for the proton density and its robust convergence with respect to . The results for are nearly indistinguishable in Fig. 1 (a). We performed similar calculations for MeV and found also robust convergence. For simplicity we show in Fig. 1 (a) only the results of MeV at . We find that the results of and MeV at are in good agreement with each other.
Based on the point-proton density distribution shown in Fig. 1 (a), we can obtain the elastic form factor straightforwardly as we mention in Sec. II. In Fig. 1 (b) we present the elastic form factor of the 4He ground state for with the same MeV and for with MeV. As we explain in Eq. (3), the elastic form factor is unity at fm-2. The NCSM results show good convergence with respect to . The converged results are in good agreement with the experimental data, which is not surprising, since the Daejeon16 interaction gives the 4He point-proton radius of fm, which is very close to the experimental value of 1.484(5) fm. Our results are also in good agreement with the EIHH calculations with the Argonne V18 (AV18) potential plus the Urbana IX (UIX) 3 force and Chiral interactions ( interaction at N3LO and 3 interaction at N2LO) in Ref. [9], which give the 4He point-proton radius of and fm, respectively. Neglecting two-body currents and further relativistic corrections, the experimental value of the point-proton radius is obtained using the experimental charge radius fm [42] via the following expression [43]:
where fm [23] is the proton finite size, fm2 [44] is the neutron finite size correction, is the mass of the proton, and the last term is a relativistic correction.
We noticed that the point-proton density and the elastic form factor have been calculated with the NCSM using the Daejeon16 interaction at in Ref. [45]. The proton form factor adopted in Ref. [45], which corresponds to the proton charge radius of fm, was taken from Ref. [46]. Our results are consistent with those in Ref. [45]. We illustrate the convergence of the these two quantities as the basis size increases.
The binding energy of 4He obtained with the Daejeon16 interaction is MeV [20]. The energy of the state is MeV which is obtained by a simple -point exponential extrapolation [21, 47] using the calculated results at and and averaged over the best convergence range MeV. The uncertainty is evaluated as half of the difference between the maximum and minimum extrapolated values over the above range. The corresponding excitation energy of the state MeV is in good agreement with the experimental value MeV.
The robust convergence and good description of the experimental data demonstrated in Fig. 1 support the reliability of our calculations of the monopole transition form factor of 4He.
In Fig. 2 we show the calculated monopole form factor of the transition in 4He at with , and MeV. We observe in Fig. 2 that the convergence of the transition monopole form factor is challenging for and MeV. The results of MeV show the best convergence. In particular, the results of MeV at and are nearly indistinguishable. The results with and MeV tend towards the converged results obtained with MeV as increases though both of these two cases do not reach apparent convergence up through [see panels (a) and (c) in Fig. 2].
In Fig. 3 we compare our results with the MAMI experimental data. We plot the NCSM results of MeV at . We also present in Fig. 3 the results of the state-of-the-art chiral interactions calculated with the HH method [17] and the EIHH method [9]. The AV18+UIX results from Ref. [9] are also shown in Fig. 3 for comparison. Our result (red curve in Fig. 3) is very close to the EIHH calculation with the chiral interactions (blue band in Fig. 3), both of which are not able to explain the experimental data. In Ref. [9], the EIHH calculation takes into account the continuum effects for the state via the Lorentz integral transform method. We calculate the state with the NCSM under the continuum discretized approximation.
We noticed that, in Ref. [14, Erratum], the Daejeon16 interaction was also used to calculate the transition monopole form factor of 4He with the NCGSM-CC method, which shows quite different results from our NCSM calculation. However, one should note that the results of Ref. [14, Erratum] were obtained by tuning the parameter to fit the experimental data on the transition form factor. We do not show also in Fig. 3 the results from Refs. [7, 15] which reproduce experimental data with simplified interactions.
IV conclusions and outlook
In conclusion, we investigated the monopole form factors in 4He with the ab initio NCSM. We performed calculations with the realistic Daejeon16 interaction. The translational invariant density and the elastic form factor of the ground state are found to be well converged as the basis size increases. The elastic form factor we calculated is in good agreement with the experimental data and the EIHH calculations with the AV18 + UIX potential and chiral +3 interactions. Our calculated monopole form factor of the transition in 4He exhibits good convergence at MeV. Our results exceed the experimental data by a sizeable amount and are comparable with the results of the EIHH calculations using the chiral (N3LO) + 3 (N2LO) interactions. From a comparison with all the available theoretical results for this observable, we see that there is a strong dependence on the Hamiltonian, as already indicated in Ref. [9], but conclude that the puzzle on this observable is still not resolved. Further investigations should be devoted to the role of the continuum in the resonance.
Acknowledgments
We acknowledge helpful discussions with Ulf-G. Meißner, Shihang Shen, Michele Viviani, Evgeny Epelbaum, Yong-Hui Lin, Pieter Maris and Nicolas Michel. A portion of the computational resources are provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC awards NP-ERCAP0020944, NP-ERCAP0023866 and NP-ERCAP0028672. This work is partially supported by new faculty startup funding by the Institute of Modern Physics, Chinese Academy of Sciences, by Key Research Program of Frontier Sciences, Chinese Academy of Sciences, Grant No. ZDB-SLY-7020, by the Natural Science Foundation of Gansu Province, China, Grant Nos. 20JR10RA067 and 23JRRA675, by the Foundation for Key Talents of Gansu Province, by the Central Funds Guiding the Local Science and Technology Development of Gansu Province, Grant No. 22ZY1QA006, by international partnership program of the Chinese Academy of Sciences, Grant No. 016GJHZ2022103FN, by National Natural Science Foundation of China, Grant No. 12375143, by National Key R&D Program of China, Grant No. 2023YFA1606903 and by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB34000000. P. Yin and B. Zhou are supported by Shanghai Research Center for Theoretical Nuclear Physics, NSFC and Fudan University, Shanghai 200438, China and the National Natural Science Foundation of China under Grant No.12147101. A. M. Shirokov is thankful to the Chinese Academy of Sciences President’s International Fellowship Initiative Program (Grant No. 2023VMA0013) which supported his visits to Lanzhou where a part of this work was performed and acknowledges the hospitality of Chinese colleagues during these visits. S.B. is supported by the Deutsche Forschungsgemeinschaft through the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1, Project ID 390831469) and through the Collaborative Research Center “Hadron and Nuclei as discovery tools” (Project ID 514321794). H. Li is supported by CSC scholarship. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under Award Nos. DE-SC0023495 and DE-SC0023692. A portion of the computational resources were also provided by Gansu Computing Center and Gansu Advanced Computing Center.